OpenStudy (anonymous):

how should i put this in my calculator to get 6 @Jhannybean it says square root (-3)^2+(3sqrt3)^2...... all of these numbers r under the sqrt sign.

OpenStudy (jhannybean):

So it's written like \[\large \sqrt{(-3)^2}+ 3(\sqrt{3})^2\]?

OpenStudy (anonymous):

aller the square root like im finding the ra

OpenStudy (jhannybean):

\[\sqrt{\large \sqrt{(-3)^2}+ 3(\sqrt{3})^2}\] like that? that'd be difficult....i suppose.

OpenStudy (jhannybean):

Ive got to head out,i asked someone to come here and help you :)

jimthompson5910 (jim_thompson5910):

draw out the problem please csmith0355

OpenStudy (anonymous):